OpenStudy (anonymous):

Rationalize the denominator. Assume that all expressions under radicals represent positive numbers. √3/√7+9

OpenStudy (kropot72):

Multipying numerator and denominator by sqrt.7 - 9 gives:\[\frac{\sqrt{3}(\sqrt{7}-9)}{(\sqrt{7}+9)(\sqrt{7}-9)}=\frac{\sqrt{21}-9\sqrt{3}}{-74}\]

OpenStudy (anonymous):

awesome thanks

OpenStudy (kropot72):

You're welcome :)